Suppose a firm finds that the marginal product of capital is 60 and the marginal product of labor is 20. If the price of capital is $6 and the price of labor is $2.50, describe how the firm should adjust its mix of capital and labor? What will be the result?
1
Expert's answer
2020-09-24T11:03:33-0400

solution

given that

marginal product of capital(MPK)=60

marginal product of labor (MPL)=20

wage rate(w)=2.50$

price of capital(r)=6$


for a firm employing labor and capital as input then best combination occur at


"\frac{MPL}{W}=\frac{MPK}{r}"


so

"\frac{MPL}{W}=\frac{20}{2.50}\\\\\\"


"\frac{MPL}{W}=8"


and


"\frac{MPK}{r}=\frac{60}{6
}"


"\frac{MPK}{r}=10"


so we can see that


"\frac{MPL}{W} <\frac{MPK}{r}\\"


"8<10"




so to attain the best input mix , the firm should hire more labor and hire lass capital so that marginal product of labor increase and marginal product of capital reduces until the equilibrium

is attained.